### 03 Feb 2024 v2

# GPT-4V with Emotion: A Zero-shot Benchmark for Generalized Emotion Recognition

Abstract
Recently, GPT-4 with Vision (GPT-4V) has demonstrated remarkable visual capabilities across various tasks, but its performance in emotion recognition has not been fully evaluated. To bridge this gap, we present the quantitative evaluation results of GPT-4V on 19 benchmark datasets covering 5 tasks: visual sentiment analysis, micro-expression recognition, facial emotion recognition, dynamic facial emotion recognition, and multimodal emotion recognition. This paper collectively refers to these tasks as “Generalized Emotion Recognition (GER)”. Through experimental analysis, we observe that GPT-4V generally outperforms supervised systems in visual sentiment analysis, highlighting its powerful visual understanding capabilities. Meanwhile, GPT-4V shows the ability to integrate multimodal clues and exploit temporal information, which is also critical for emotion recognition. Despite these achievements, GPT-4V is primarily tailored for general-purpose domains, which cannot recognize micro-expressions that require specialized knowledge. To the best of our knowledge, this paper provides the first quantitative assessment of GPT-4V for the GER tasks, offering valuable insights to researchers in this field. It can also serve as a zero-shot benchmark for subsequent research. Our code and evaluation results are available at: https://github.com/zeroQiaoba/gpt4v-emotion.
keywords:
Generalized Emotion Recognition (GER); GPT-4 with Vision (GPT-4V); Zero-shot Benchmark; Multimodal Fusion; Temporal Modeling.
[…]
[label1] organization=Institute of Automation, Chinese Academy of Sciences, city=Beijing, country=China
[…]
[label2] organization=School of Artificial Intelligence, University of Chinese Academy of Sciences, city=Beijing, country=China
[…]
[label3] organization=Software Engineering Institute, Peking University, city=Beijing, country=China
[…]
[label4] organization=Department of Automation, Tsinghua University, city=Beijing, country=China
[…]
[label5] organization=Beijing National Research Center for Information Science and Technology, Tsinghua University, city=Beijing, country=China
1 Introduction
Emotion has attracted increasing attention from researchers due to its importance in human-computer interaction cowie2001emotion ; brave2007emotion . Current research on emotion recognition primarily centers on two aspects: one aims to identify the emotions evoked by stimuli, essentially predicting how viewers might feel after seeing these stimuli you2017visual ; zhao2021affective ; the other aims to analyze human emotions in images li2019blended ; li2020deep or videos lian2022smin ; lian2024merbench . This paper collectively refers to these tasks as “Generalized Emotion Recognition (GER)”.
Emotion can be conveyed through various modalities, including text, audio, and video picard2000affective . Among them, visual information (such as colorfulness, brightness, facial expression, human action, etc) contains rich emotion-related content yang2023emoset and plays a crucial role in GER tasks. To enhance visual understanding capabilities, researchers have proposed various algorithms and achieved noteworthy progress poria2017review . With the development of deep learning, current research on GER has shifted from handicrafted features to deep neural networks. Recently, GPT-4V openai2023gpt4v has showcased impressive visual understanding capabilities across various tasks. This leads to a question: can GPT-4V solve the GER problem to some extent? If so, what is the future direction following the emergence of GPT-4V?
In September 2023, GPT-4V was integrated into ChatGPT openai2022chatgpt , triggering a series of user reports aimed at investigating its visual capabilities yang2023dawn ; lyu2023gpt . However, these reports generally involve a limited number of samples per task and can only provide qualitative insights into GPT-4V. Although OpenAI released the API in November 2023, it was initially limited to 100 requests per day. Therefore, it remains difficult to evaluate GPT-4V against state-of-the-art systems on benchmark datasets. Recently, OpenAI has increased the daily limit, allowing us to conduct more comprehensive evaluations.
In this paper, we provide quantitative evaluation results of GPT-4V on GER tasks, covering visual sentiment analysis ortis2020survey , micro-expression recognition li2022deep , facial emotion recognition li2020deep , dynamic facial emotion recognition wang2022ferv39k , and multimodal emotion recognition lian2024merbench . Figure 1 shows the overall results of GPT-4V. We also report results of random guessing (a heuristic baseline that randomly selects labels from candidate categories) and supervised systems. To fairly compare different methods, we conduct experiments on benchmark datasets and use the same evaluation metric. As shown in Figure 1, overall, GPT-4V outperforms random guessing but still lags behind supervised systems. To figure out the reason, we conduct a comprehensive analysis of GPT-4V’s multi-faceted performance, including multimodal fusion ability, temporal modeling capability, color space robustness, prediction consistency, etc (see Section 5). We hope this paper can inspire subsequent researchers and shed light on tasks that GPT-4V can effectively address and those that require further exploration. The main contribution of this paper can be summarized as follows:
  • 1.
    To the best of our knowledge, this is the first work to quantitatively evaluate the performance of GPT-4V in emotion recognition. Additionally, this paper can serve as a zero-shot benchmark for subsequent research.
  • 2.
    Our evaluation results demonstrate that GPT-4V focuses on general-purpose domains. For micro-emotion recognition that needs specialized knowledge, it performs even worse than random guessing in some datasets. Meanwhile, we observe that GPT-4V can integrate multimodal clues and capture temporal information during inference.
  • 3.
    This paper can offer guidance for future directions following the emergence of GPT-4V. We open-source the code and encourage subsequent researchers to broaden the evaluation scope by including more tasks and datasets.
The remainder of this paper is organized as follows: In Section 2, we briefly review recent works. In Sections 3 and 4, we provide a comprehensive overview of GER tasks and datasets, along with a detailed description of the GPT-4V calling strategy. In Section 5, we report results and conduct an in-depth analysis of GPT-4V’s multi-faceted capabilities. Finally, we conclude this paper and discuss future work in Section 6.
Figure 1: Performance of different methods on GER. “Random” is a heuristic baseline that randomly selects labels from candidate categories.
2 Related Works
2.1 Generalized Emotion Recognition
GER aims to identify the emotions of individuals or to recognize the emotions that individuals might evoke from stimuli. It covers a variety of tasks, and this paper focuses on five representative GER tasks. In this section, we detail the similarities and distinctions between these tasks.
Visual sentiment analysis ortis2020survey investigates the impact of image stimuli on people’s emotions, where the image content is not necessarily human-centric. In contrast, facial emotion recognition li2020deep aims to identify emotions in human-centric images, particularly focusing on observable macro-expressions. In addition to macro-expressions, there is another kind of expression called micro-expressions, which are characterized by short duration, low intensity, and sparse facial action units. The task of identifying such expressions is called micro-expression recognition li2022deep . All of the aforementioned tasks identify emotions from static images. Dynamic facial emotion recognition wang2022ferv39k extends the analysis from static images to image sequences, requiring the use of temporal information. Moreover, emotions can be conveyed through modalities other than visuals, such as text and audio. Therefore, multimodal emotion recognition lian2024merbench is introduced to integrate multimodal information for a more comprehensive understanding of emotions.
2.2 Multimodal Large Language Model
Recently, large language models (LLMs) with massive parameters have attracted increasing attention due to their unprecedented performance penedo2023refinedweb ; touvron2023llama ; workshop2022bloom . In GER tasks, visual information plays an important role in emotion recognition, but LLMs only support lexical inputs. Therefore, this section focuses on multimodal LLMs that support visual inputs.
Among all multimodal LLMs, MiniGPT-4 zhu2023minigpt and LLaVA liu2023visual are two representative models. They enhance the visual capability by aligning the frozen visual encoder with LLMs using a simple projection layer., which preserves most of the information encoded in pre-trained models and achieves good performance in visual understanding. This approach is further extended by other works and applies to support more modalities, such as video li2023videochat , audio zhang2023speechgpt , and depth images su2023pandagpt . Currently, GPT-4V openai2023gpt4v is the state-of-the-art multimodal LLM for various tasks alpaca_eval ; lu2024mathvista . However, its performance in emotion recognition has not been fully studied. Hence, this paper evaluates the performance of GPT-4V on GER, providing a valuable zero-shot benchmark for other multimodal LLMs.
2.3 Evaluation on GPT-4V
With the emergence of GPT-4V, a large number of user reports have surfaced to evaluate its visual capabilities lin2023mm ; zhou2023exploring ; wu2023can . Among them, Yang et al. yang2023dawn conducted the most comprehensive assessment, covering coding, captioning, and reasoning abilities. However, they primarily performed qualitative analysis on limited samples and did not disclose the performance gap between GPT-4V and supervised systems on benchmark datasets. To bridge this gap, there have been efforts to quantitatively analyze GPT-4V’s mathematical reasoning lu2024mathvista and visual abilities wu2023gpt4vis , but none have specifically explored its emotional understanding capabilities. In this paper, we provide the first quantitative evaluation of GPT-4V on GER tasks. It is worth noting that, unlike other tasks, emotions are related to humans. However, human identity is treated as sensitive information in GPT-4V, which raises new requirements for the calling strategy. More details can be found in Section 4.
Table 1: Statistics for different datasets. For each dataset, we evaluate performance on the official test set and select the most common evaluation metric. In this table, ACC, WAR, and WAF represent accuracy, weighted average recall, and weighted average F-score, respectively.
Task (Modality) Dataset # Test samples Metric Labels
Visual Sentiment Analysis (Image) Twitter I you2015robust 1,269 ACC positive, negative
Twitter II borth2013large 603 ACC positive, negative
Abstract you2016building 228 ACC amusement, sadness, anger, fear, disgust, awe, content, excitement
ArtPhoto you2016building 806 ACC amusement, sadness, anger, fear, disgust, awe, content, excitement
Micro-expression Rec. (Image) CASME yan2013casme 195 ACC happiness, sadness, fear, surprise, disgust, repression, contempt, tense
CASME II yan2014casme 247 ACC happiness, sadness, fear, surprise, disgust, repression, others
SAMM davison2016samm 159 ACC happiness, sadness, fear, surprise, disgust, contempt, anger, others
Facial Emotion Rec. (Image) CK+ lucey2010extended 981 ACC happiness, sadness, anger, fear, disgust, surprise, contempt
SFEW 2.0 dhall2015video 436 ACC happiness, sadness, anger, fear, disgust, surprise, neutral
RAF-DB li2017reliable 3,068 ACC happiness, sadness, anger, fear, disgust, surprise, neutral
FERPlus barsoum2016training 3,589 ACC happiness, sadness, anger, fear, disgust, surprise, neutral, contempt
AffectNet mollahosseini2017affectnet 4,000 ACC happiness, sadness, anger, fear, disgust, surprise, neutral, contempt
Dynamic Facial Emotion Rec. (Video) DFEW (fd1) jiang2020dfew 2,341 WAR happiness, sadness, anger, fear, disgust, surprise, neutral
FERV39k wang2022ferv39k 7,847 WAR happiness, sadness, anger, fear, disgust, surprise, neutral
RAVDESS livingstone2018ryerson 1,440 WAR happiness, sadness, anger, fear, disgust, surprise, neutral, calm
eNTERFACE05 martin2006enterface 1,287 WAR happiness, sadness, anger, fear, disgust, surprise
Multimodal Emotion Rec. (Video, Audio, Text) CMU-MOSI zadeh2017tensor 686 WAF sentiment intensity
CH-SIMS yu2020ch 457 WAF sentiment intensity
MER-MULTI lian2023mer 411 WAF happiness, sadness, anger, surprise, neutral, worry
Figure 2: Examples of different datasets. We pixelate faces due to the sensitive nature of human identity.
3 Task Description
In this paper, we conduct a comprehensive evaluation across 19 datasets covering 5 tasks. Table 1 summarizes statistics of different datasets. To ensure a fair comparison with supervised systems, we evaluate performance on the official test set and select the most common evaluation metric. Figure 2 shows examples of different datasets, demonstrating the diversity of dataset types. Some datasets are in-the-wild (such as AffectNet), while others are laboratory-controlled (such as CASME and CK+). Meanwhile, there is a disparity in color spaces across distinct datasets. Some datasets use grayscale images (such as CK+), while others use RGB images (such as CASME and AffectNet). This section describes each task and dataset in detail.
Visual Sentiment Analysis aims to identify the emotions that individuals might evoke from images. We focus on four datasets: Twitter I you2015robust , Twitter II borth2013large , ArtPhoto you2016building , and Abstract you2016building . Specifically, Twitter I you2015robust and Twitter II borth2013large are collected from social websites. Of the two, Twitter I provides raw annotations from 5 Amazon Mechanical Turk workers. Consistent with previous works, we employ three label processing methods, resulting in Twitter I(3), Twitter I(4), and Twitter I(5). Taking Twitter I(4) as an example, it contains all samples where at least 4 workers assign the same label. Different from Twitter I and Twitter II, ArtPhoto you2016building contains artistic photographs from a photo-sharing site, while Abstract you2016building consists of peer-rated abstract paintings. Both datasets are annotated into eight categories. In line with previous works, we reassign these labels into positive and negative classes and report results for the negative/positive classification task.
Facial Emotion Recognition chooses five benchmark datasets: CK+ lucey2010extended , FERPlus barsoum2016training , SFEW 2.0 dhall2015video , RAF-DB li2017reliable , and AffectNet mollahosseini2017affectnet . The first two datasets contain grayscale images, while the last three datasets contain RGB images. Specifically, CK+ lucey2010extended includes 593 video sequences from 123 subjects. Following previous work jiang2022disentangling , we extract the last three frames of each sequence to construct the dataset. FERPlus barsoum2016training is an extension of FER2013 goodfellow2013challenges , where each sample is relabeled by ten annotators. SFEW 2.0 dhall2015video extracts keyframes from movie clips, encompassing diverse head poses, occlusions, and illuminations. RAF-DB li2017reliable contains thousands of samples with basic and compound expressions, and this paper focuses on basic emotions. AffectNet mollahosseini2017affectnet has 8 labels and each label contains 500 samples for evaluation.
Table 2: Prompt template for different tasks. Please replace “Candidate labels” in the following prompts into labels of each dataset. This paper uses batch-wise calling strategy to handle request limits. By default, we set the batch size to 20 for image-level inputs and 6 for video-level inputs.
Task
Prompt
Visual Sentiment Analysis
Please play the role of an emotion recognition expert. We provide 20 images. Please recognize sentiments evoked by these images (i.e., guess how viewer might emotionally feel after seeing these images.) If there is a person in the image, ignore that person’s identity. For each image, please sort the provided categories from high to low according to the similarity with the image. Here are the optional categories: {Candidate labels}. The output format should be {’name’:, ’result’:} for each image.
Micro-expression Rec.
Please play the role of a micro-expression recognition expert. We provide 20 images. For each image, please sort the provided categories from high to low according to the similarity with the input image. The expression may not be obvious, please pay attention to the details of the face. Here are the optional categories: {Candidate labels}. Please ignore the speaker’s identity and focus on the facial expression. The output format should be {’name’:, ’result’:} for each image.
Facial Emotion Rec.
Please play the role of a facial expression classification expert. We provide 20 images. For each image, please sort the provided categories from high to low according to the top 5 similarity with the input image. Here are the optional categories: {Candidate labels}. Please ignore the speaker’s identity and focus on the facial expression. The output format should be {’name’:, ’result’:} for each image.
Dynamic Facial Emotion Rec.
Please play the role of a video expression classification expert. We provide 6 videos, each with 3 temporally uniformly sampled frames. For each video, please sort the provided categories from high to low according to the top 5 similarity with the input video. Here are the optional categories: {Candidate labels}. Please ignore the speaker’s identity and focus on the facial expression. The output format should be {’name’:, ’result’:} for each video.
Multimodal Emotion Rec.
Please play the role of a video expression classification expert. We provide 6 videos, each with the speaker’s content and 3 temporally uniformly sampled frames. Please ignore the speaker’s identity and focus on their emotions. For each video, please sort the provided categories from high to low according to the top 5 similarity with the input video. Here are the optional categories: {Candidate labels}. Please ignore the speaker’s identity and focus on their emotions. The output format should be {’name’:, ’result’:} for each video.
Micro-expression Recognition aims to identify subtle changes in human faces. Early works have proved the feasibility of recognizing micro-expressions from an apex frame li2020joint . Therefore, we use the apex frame in our evaluation. Following previous works, we adopt the same label processing method and concentrate on major emotions. Specifically, CASME yan2013casme comprises 195 samples across 8 categories, and we focus on the performance of four main labels: tense, disgust, repression, and surprise. CASME II yan2014casme includes 247 samples collected from 26 subjects, and we focus on five major labels: happiness, surprise, disgust, repression, and others. SAMM davison2016samm consists of 159 samples, and our evaluation is limited to labels with more than 10 samples: anger, contempt, happiness, surprise, others.
Dynamic Facial Emotion Recognition focuses on more challenging image sequences. In this task, we conduct experiments on four benchmark datasets: FERV39k wang2022ferv39k , RAVDESS livingstone2018ryerson , eNTERFACE05 martin2006enterface , and DFEW jiang2020dfew . For the first three datasets, we use the official training/validation/test splits and evaluate the performance on the official test set. DFEW jiang2020dfew contains 5 folds with a total of 11,697 samples. To reduce the evaluation cost, we only report the results of fold 1 (fd1), consistent with previous works wang2022dpcnet ; chen2023static .
Multimodal Emotion Recognition aims to integrate various modalities (including audio, video, and text) to identify emotions. For this task, we select three benchmark datasets: CH-SIMS zadeh2017tensor , CMU-MOSI yu2020ch , and MER-MULTI lian2023mer . The first two datasets offer sentiment intensity scores for each sample. In our evaluation, we concentrate on the negative/positive classification task, where positive and negative classes are assigned for $<0$ and $>0$ scores, respectively. MER-MULTI is a subset of the MER2023 dataset lian2023mer , providing both discrete and dimensional labels for each sample. In this paper, we focus on the recognition performance of discrete emotions.
4 GPT-4V Calling Strategy
This paper evaluates the performance of the latest GPT-4V API, “gpt-4-vision-preview”. GER tasks involve diverse modalities, such as image, text, video, and audio. However, the current version of GPT-4V has limitations and only supports image and text inputs. To process video data, we sample the video and convert it into multiple images. For audio data, we try converting it to a mel-spectrogram, but GPT-4V fails to generate proper responses for this input. Therefore, the primary focus of this paper revolves around image, text, and video. To meet the request limits of API and reduce rejected cases caused by security checks, we propose batch-wise calling and recursive calling strategies. More details are provided below.
Input: Dataset $\mathcal{D}=\{x_{i}\}_{i=1}^{N}$ with ${N}$ samples, the maximum number of calls $T_{\max}$.
Output: Prediction results for all samples.
1 Def Main($\mathcal{D}$):
2       Split $\mathcal{D}$ into batches $\mathcal{D}=\{\mathcal{B}_{i}\}_{i=1}^{N_{b}}$; for $i=1,\cdots,N_{b}$ do
3             Func($\mathcal{B}_{i}$);
4       end for
5      return
6
7
8 Def Func($\mathcal{B}_{i}$):
9       for $t=1,\cdots,T_{\max}$ do
10             Call GPT-4V for $\mathcal{B}_{i}$ and get outputs $\mathcal{O}_{i}$; if passed checks then
11                   Store the outputs $\mathcal{O}_{i}$; return
12             end if
13
14       end for
15      if $|\mathcal{B}_{i}|\geq 2$ then
16             Divide $\mathcal{B}_{i}$ into two equal parts: $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$; Func($\mathcal{C}_{1}$); Func($\mathcal{C}_{2}$);
17       end if
18      return
19
Algorithm 1 GPT-4V Calling Strategy
4.1 Batch-wise Calling Strategy
The current GPT-4V API has three limitations on requests: tokens per minute (TPM), requests per minute (RPM), and requests per day (RPD), which introduces additional requirements to our prompt design. To meet the constraints of RPM and RPD, we follow previous work wu2023gpt4vis and adopt batch-wise inputs. That is, we feed multiple samples into GPT-4V and request it to generate the response for each sample. However, a large batch size may result in the total number of tokens exceeding the TPM limit. Additionally, it increases task difficulty and may lead to incorrect outputs. For example, if we input 30 samples in a batch, we might only receive 28 predictions. Therefore, we set the batch size to 20 for image-level inputs and 6 for video-level inputs, aiming to satisfy three API limitations (i.e., TPM, RPM, and RPD) simultaneously. Prompts for different tasks can be found in Table 2.
4.2 Recursive Calling Strategy
During the evaluation, GER tasks tend to trigger security checks in GPT-4V. This is attributed to the inclusion of visual sentiment analysis and human emotion recognition. The former task contains violent and bloody images. In the latter task, human identities are also considered as sensitive information. To reduce these errors, we require GPT-4V to ignore speaker identity in the prompt (see Table 2), but it still triggers security errors. Interestingly, these errors sometimes occur randomly. For example, although all images are human-centric, some pass security checks while others fail. Alternatively, a sample may initially fail the check but pass upon retry. Therefore, we make multiple calls to the same batch to reduce rejected cases.
Meanwhile, we notice that a batch input may fail security checks, but splitting it into smaller parts can sometimes pass the checks. Therefore, for consistently failing batches, we split them into two smaller mini-batches and feed these mini-batches into GPT-4V. This operation is repeated until further splitting is no longer feasible. We call this strategy the “recursive calling strategy” and its pseudo-code is summarized in Algorithm 1.
5 Results and Discussion
In this section, we first show the performance of the heuristics baselines, supervised systems, and GPT-4V on GER tasks. Then, we reveal the multimodal fusion and temporal modeling capabilities of GPT-4V. Following that, we visualize confusion matrices and conduct an in-depth class-wise performance analysis. Next, we analyze the system stability and explore the color space robustness. Finally, we visualize the samples that are rejected in the security check and erroneously predicted.
Table 3: Performance comparison on visual sentiment analysis.
Model Twitter I (5/4/3) Twitter II Abstract ArtPhoto
SentiBank borth2013large 71.32/68.28/66.63 65.93 64.95 67.74
PAEF zhao2014exploring 72.90/69.61/67.92 77.51 70.05 67.85
DeepSentiBank chen2014deepsentibank 76.35/70.15/71.25 70.23 71.19 68.73
PCNN you2015robust 82.54/76.52/76.36 77.68 70.84 70.96
VGGNet simonyan2015very 83.44/78.67/75.49 71.79 68.86 67.61
AR (Concat) yang2018visual 88.65/85.10/81.06 80.48 76.03 74.80
Random 50.74/49.75/50.50 50.37 50.88 50.41
Majority 66.82/62.51/61.33 78.61 61.23 53.06
GPT-4V 97.81/94.63/90.71 87.95 71.81 80.40
Table 4: Performance comparison on micro-expression recognition.
Model CASME CASME II SAMM
LBP-SIP wang2015lbp 36.84 46.56 36.76
STRCN-A xia2019spatiotemporal 40.93 45.26 32.85
STRCN-G xia2019spatiotemporal 59.65 63.37 53.48
VGGMag li2018can 60.23 63.21 36.00
LGCcon li2020joint 60.82 65.02 40.90
TSCNN song2019recognizing 73.88 80.97 71.76
Random 24.09 18.33 20.37
Majority 40.34 40.17 42.22
GPT-4V 36.93 14.64 17.04
Table 5: Performance comparison on facial emotion recognition.
Model CK+ Model SFEW 2.0 Model RAF-DB Model FERPlus Model AffectNet
Cross-VAE wu2020cross 94.96 IACNN meng2017identity 50.98 SCN wang2020suppressing 87.03 RANwang2020region 89.16 SCN wang2020suppressing 60.23
IA-gen yang2018identity 96.57 DLP-CNN li2017reliable 51.05 EfficientFace zhao2021robust 88.36 SCN wang2020suppressing 89.39 MA-Net zhao2021learning 60.29
ADFL bai2019disentangled 98.17 TDTLN yan2019cross 53.10 MA-Net zhao2021learning 88.42 EAC zhang2022learn 89.64 ARM shi2021learning 61.33
IDFERM liu2019hard 98.35 DAN wen2023distract 53.18 TransFER xue2021transfer 90.91 KTNli2021adaptively 90.49 DAN wen2023distract 62.09
TER-GAN ali2021facial 98.47 RAN wang2020region 54.19 POSTER zheng2023poster 92.05 TransFERxue2021transfer 90.83 POSTER zheng2023poster 63.34
IPD-FER jiang2022disentangling 98.65 FaceCaps wu2021facecaps 58.50 POSTER++ mao2023poster 92.21 POSTER zheng2023poster 91.62 POSTER++ mao2023poster 63.77
Random 14.61 Random 13.98 Random 14.57 Random 12.61 Random 12.73
Majority 25.38 Majority 19.77 Majority 38.64 Majority 35.75 Majority 12.50
GPT-4V 69.72 GPT-4V 57.24 GPT-4V 75.81 GPT-4V 64.25 GPT-4V 42.77
Table 6: Performance comparison on dynamic facial emotion recognition. Since GPT-4V does not support video input directly, we uniformly sample frames from the videos and sequentially feed these frames into GPT-4V. In this table, we test 2-frame (2-frm) and 3-frame (3-frm) sampling manners.
Model DFEW Model FERV39k Model RAVDESS Model eNTERFACE05
C3D tran2015learning 56.37 M3DFEL wang2023rethinking 47.67 AV-LSTM ghaleb2019multimodal 65.80 STA-FER pan2019deep 42.98
ResNet50-LSTM wang2022dpcnet 63.32 STT ma2022spatio 48.11 MCBP su2020msaf 71.32 EC-LSTM miyoshi2021enhanced 49.26
DPCNet wang2022dpcnet 65.78 IAL li2023intensity 48.54 MSAF su2020msaf 74.86 FAN meng2019frame 51.44
SVFAP sun2023svfap 74.81 MAE-DFER sun2023mae 52.07 SVFAP sun2023svfap 75.01 Graph-Tran zhao2022spatial 54.62
MAE-DFER sun2023mae 74.88 SVFAP sun2023svfap 52.29 MAE-DFER sun2023mae 75.56 SVFAP sun2023svfap 60.54
S2D chen2023static 76.16 S2D chen2023static 52.56 CFN-SR fu2021cross 75.76 MAE-DFER sun2023mae 61.64
Random 14.43 Random 14.41 Random 11.86 Random 16.87
Majority 22.81 Majority 24.79 Majority 13.33 Majority 16.71
GPT-4V (2-frm) 43.80 GPT-4V (2-frm) 34.29 GPT-4V (2-frm) 22.29 GPT-4V (2-frm) 23.78
GPT-4V (3-frm) 54.80 GPT-4V (3-frm) 38.72 GPT-4V (3-frm) 34.31 GPT-4V (3-frm) 33.28
Table 7: Performance comparison on multimodal emotion recognition. These results come from MERBench lian2024merbench , which reproduces different algorithms using the same feature set to ensure a fair comparison.
Model CMU-MOSI CH-SIMS MER-MULTI
MFM tsai2018learning 78.34 87.14 70.48
MISA hazarika2020misa 79.08 89.82 82.42
MFN zadeh2018memory 77.78 87.00 77.40
MMIM han2021improving 79.38 88.68 81.01
TFN zadeh2017tensor 79.63 90.56 82.52
MulT tsai2019multimodal 81.41 91.07 82.94
Random 51.33 51.23 17.87
Majority 42.37 50.47 10.40
GPT-4V (3-frm) 80.43 81.24 65.39
5.1 Main Results
In this section, we report the performance of different methods on five GER tasks. In Tables 3$\sim$7, we include two additional heuristic baselines: random guessing and majority guessing. For the former, we randomly select a label from the candidate categories as the prediction. For the latter, we choose the majority label as the prediction. For both baselines, we run experiments 10 times and report the average results.
Table 3 shows the results of visual sentiment analysis. From this table, we observe that GPT-4V outperforms supervised systems on most datasets. This superior performance can be attributed to its strong understanding of image content, coupled with its reasoning capabilities, allowing GPT-4V to accurately infer the emotional states evoked by the images. But for micro-emotion recognition (see Table 4), GPT-4V exhibits poor performance, even worse than heuristic baselines. These results suggest that GPT-4V is designed for general-purpose domains (i.e., emotions that can be perceived by normal people). It is not well-suited for tasks that require professional knowledge.
Tables 5$\sim$7 show the performance of facial emotion recognition, dynamic facial emotion recognition, and multimodal emotion recognition, respectively. The latter two tasks identify emotions in videos. Since GPT-4V does not support video input, we uniformly sample frames from the video and then feed these frames into GPT-4V sequentially. To reduce GPT-4V call costs, we sample up to three frames per video. Despite a performance gap still existing between GPT-4V and supervised systems, it is worth noting that GPT-4V significantly outperforms the heuristic baselines. These results demonstrate its potential in emotion recognition. In Table 7, GPT-4V performs well on CMU-MOSI but relatively poorly on MER-MULTI. This discrepancy can be attributed to the fact that acoustic information is more important in MER-MULTI compared to CMU-MOSI lian2024merbench . Since GPT-4V does not support audio input, MER-MULTI loses more information, leading to limited performance.
5.2 Multimodal Understanding Ability
In this section, we evaluate the multimodal understanding capabilities of GPT-4V. Among all tasks, only multimodal emotion recognition provides multimodal information. Therefore, we conduct experiments on this task. Table 8 reports both unimodal and multimodal results. For CH-SIMS and MER-MULTI, we observe that multimodal results outperform unimodal results, showcasing GPT-4V’s ability to integrate and leverage multi-modalities. But for CMU-MOSI, we note a slight decrease in multimodal results compared to unimodal results. This can be attributed to CMU-MOSI primarily relying on lexical information to convey emotions, and the incorporation of visual information may introduce interference lian2024merbench .
5.3 Temporal Understanding Ability
To reduce GPT-4V call costs, this paper samples up to three frames per video. In this section, we further reveal the impact of sampling numbers. In Table 6, the performance improves when we increase the sampling number from two to three, indicating GPT-4V’s temporal understanding ability. Additionally, it is worth noting that despite setting the sampling number to three, some key frames may still be ignored. Therefore, there is a possibility for further improvement by sampling more frames, and we leave this discussion for subsequent researchers.
Table 8: Analysis of multimodal fusion ability. We provide both unimodal and multimodal results for the multimodal emotion recognition task.
Modality CMU-MOSI CH-SIMS MER-MULTI
Text 82.32 70.07 34.57
Video 51.17 76.13 46.23
Multi 80.43 81.24 65.39
(a) Twitter I(5)
(b) Twitter I(4)
(c) Twitter I(3)
(d) Twitter II
(e) Abstract
(f) ArtPhoto
(g) CASME
(h) CASME II
(i) SAMM
(j) CK+
(k) SFEW 2.0
(l) FERPlus
(m) RAF-DB
(n) AffectNet
(o) DFEW
(p) FERV39k
(q) RAVDESS
(r) eNTERFACE05
(s) CMU-MOSI
(t) CH-SIMS
(u) MER-MULTI
Figure 3: Confusion matrices for all evaluation datasets. Different colors represent distinct tasks.
5.4 Confusion Matrix
In Figure 3, we visualize confusion matrices and analyze the performance of GPT-4V across various emotion categories.
For visual sentiment analysis, GPT-4V generally performs well in both positive and negative classes, showcasing its robust visual understanding capabilities. However, it exhibits lower performance on the Abstract dataset compared to other datasets. This may be attributed to the fact that GPT-4V is mainly trained on natural images. The domain gap between abstract and natural images leads to increased difficulty in emotion recognition.
For micro-expression recognition, GPT-4V prefers tense and repression. Compared to other labels (e.g., surprise and disgust), tense and repression are less noticeable and closer to neutral. These results indicate that GPT-4V tends to view micro-expressions as neutral, causing it to perform poorly on this task.
For facial expression recognition, GPT-4V tends to exhibit lower accuracy on contempt, disgust, and fear, while achieving higher accuracy on happiness. This phenomenon is related to the clarity of emotion definition. Since contempt, disgust, and fear have less consensus between annotations, they are inherently more difficult to recognize mollahosseini2017affectnet ; picard2001toward .
For dynamic facial emotion recognition, most samples tend to be predicted as neutral, which is more obvious in the FERV39 dataset. We attribute this phenomenon to our sampling operation. For each video, we sample up to three frames. If there is no emotion-related information within these frames, GPT-4V will predict the emotion as neutral.
For multimodal emotion recognition, GPT-4V performs well on worry and happiness. For worry, it has a strong reliance on lexical information. Differently, happiness is mainly expressed through facial expressions lian2023mer . To achieve good performance, GPT-4V should use different predominant modalities for distinct emotions, showcasing its multimodal fusion capabilities.
(a) identical predictions
(b) test accuracy
Figure 4: System stability test. We run GPT-4V 10 times, calculate the frequency with identical predictions, and report the test accuracy for different runs.
5.5 System Stability
In this section, we evaluate the system stability of GPT-4V. Experiments are conducted on SFEW 2.0, using GPT-4V to predict each sample 10 times. Figure 4 calculates the frequency with identical predictions and reports the test accuracy for different runs. As shown in Figure 4a, although most samples exhibit identical predictions across 10 runs, there are some samples with different predictions, leading to fluctuations in the test accuracy. In Figure 4b, we observe that the maximum performance gap between the best and worst runs is 4.60%. Hence, GPT-4V exhibits some instability. We recommend that subsequent researchers evaluate GPT-4V in multiple runs and consider the majority predictions as the final results.
5.6 Robustness to Color Space
As illustrated in Figure 2, there are differences in the color spaces across distinct datasets. In this section, we evaluate the robustness of GPT-4V to color space changes. We conduct experiments on RAF-DB and convert all RGB images to grayscale. We observe that GPT-4V achieves nearly identical performance on grayscale images (74.28%) and RGB images (75.81%). These results demonstrate its color space robustness.
Figure 5 further visualizes the prediction consistency on grayscale and RGB images. We observe that the consistency rate is around 80% for most emotion categories, except for fear and disgust. We further compare Figure 5 with the confusion matrix of RAF-DB (see Figure 3m). We observe significant similarities between these images, where fear and disgust also perform poorly in emotion recognition. Therefore, a reasonable conjecture emerges: poorly-performing emotion categories in GPT-4V may exhibit greater randomness across different runs.
Figure 5: Prediction consistency on RGB and grayscale images.
5.7 Security Check
GER includes visual sentiment analysis (containing images with violent and bloody content) and human emotion recognition (where human identity is considered sensitive information in GPT-4V). Therefore, it often triggers the security check. Figure 6 shows passed and rejected examples. We observe that all rejected examples contain humans or violent scenes. But there are some cases where examples containing such information can still pass checks. This observation suggests a certain degree of instability in the current security check used by GPT-4V.
5.8 Case Study
In Figure 7, we visualize examples where GPT-4V makes incorrect predictions. More examples can be found in Figures 8$\sim$12 in the Appendix. For visual sentiment analysis (see Twitter I), we notice that GPT-4V can generate reasonable predictions. The main reason for the errors is the inherent multiple meanings of images. Take the image on the left as an example. One interpretation could emphasize that the sun rises, breaking through dark clouds and bathing the trees in sunlight, evoking positive emotions. Conversely, another explanation could emphasize that the sun is obstructed by dark clouds, triggering negative emotions. The same phenomenon can also be found in the picture on the right. On the one hand, there is a cute cat, but on the other hand, the cat is locked in a cage.
Figure 6: Passed and rejected examples by GPT-4V.
Figure 7: Error cases predicted by GPT-4V.
For micro-expression recognition, we observe that her emotion is not obvious (see CASME in Figure 7). Without specialized knowledge, it is difficult to identify her emotion as disgust and more likely to be predicted as neutral. For CASME, there are four candidate labels: tense, disgust, repression, and surprise. Among them, tense is closer to neutral. Consequently, GPT-4V incorrectly predicts her emotion as tense. In the future, we recommend that researchers provide clearer definitions of micro-expressions in the prompt design. For example, offer typical clues to distinguish these emotions. We believe this strategy can enhance the performance of GPT-4V on this task.
For facial emotion recognition (see RAF-DB), the image depicts a baby with an open mouth, indicating that she is crying. Meanwhile, she is wearing an oxygen mask, indicating that she is sick. Hence, her emotion is likely to be sad. But GPT-4V incorrectly predicts her emotion as neutral. This may be due to its inability to recognize the oxygen mask and infer her health status, indicating its limitations in understanding the details.
For dynamic facial emotion recognition (see DFEW), GPT-4V might predict the emotion as happiness based on the raised corners of the mouth. But when considering the subtitle, “cannot go to the academy, read books, and listen to music”, we can infer that he is more likely to be sad due to these restrictions. Hence, the error made by GPT-4V could be attributed to the limitations of this task, as it relies only on the visual modality.
For multimodal emotion recognition (see CH-SIMS), GPT-4V predicts the emotion as positive based on his facial expression. However, when taking into account the subtitle, “I feel like I’m falling apart”, the content of his complaint shifts his emotion towards negative. However, GPT-4V is unable to capture such emotional transitions, indicating its limitations in multimodal fusion under complex scenarios.
In summary, GPT-4V provides relatively reasonable predictions. These errors are primarily caused by the inherent multiple meanings of images or the limited capability to identify small objects. Additionally, micro-expressions are different from regular expressions, and GPT-4V encounters challenges in recognizing them without specialized knowledge. Furthermore, its multimodal fusion ability may face challenges when different modalities express distinct emotions.
6 Conclusion
This paper conducts a comprehensive evaluation of GPT-4V in GER tasks. It is worth noting that GPT-4V exhibits strong visual understanding capabilities, even surpassing supervised systems in visual sentiment analysis. However, it performs poorly in micro-expression recognition which requires specialized knowledge. This paper also showcases GPT-4V’s temporal modeling and multimodal fusion capabilities, along with its robustness to color space changes. Furthermore, we evaluate its prediction consistency and security check stability, and visualize error cases to reveal its limitations in emotion understanding. This paper can serve as a zero-shot benchmark and provide guidance for subsequent research on emotion recognition and multimodal LLMs. In the future, we will expand the evaluation scope by including more emotion-related tasks and datasets.
Acknowledgements
This work is supported by the National Natural Science Foundation of China (NSFC) (No.62201572, No.61831022, No.62276259, No.U21B2010), Beijing Municipal Science & Technology Commission, Administrative Commission of Zhongguancun Science Park No.Z211100004821013, Open Research Projects of Zhejiang Lab (NO. 2021KH0AB06), and CCF-Baidu Open Fund (No.OF2022025).