### 18 Mar 2024 v3

# GPT-4V with Emotion: A Zero-shot Benchmark for Generalized Emotion Recognition

Abstract
Recently, GPT-4 with Vision (GPT-4V) has demonstrated remarkable visual capabilities across various tasks, but its performance in emotion recognition has not been fully evaluated. To bridge this gap, we present the quantitative evaluation results of GPT-4V on 21 benchmark datasets covering 6 tasks: visual sentiment analysis, tweet sentiment analysis, micro-expression recognition, facial emotion recognition, dynamic facial emotion recognition, and multimodal emotion recognition. This paper collectively refers to these tasks as “Generalized Emotion Recognition (GER)”. Through experimental analysis, we observe that GPT-4V exhibits strong visual understanding capabilities in GER tasks. Meanwhile, GPT-4V shows the ability to integrate multimodal clues and exploit temporal information, which is also critical for emotion recognition. However, it’s worth noting that GPT-4V is primarily designed for general domains and cannot recognize micro-expressions that require specialized knowledge. To the best of our knowledge, this paper provides the first quantitative assessment of GPT-4V for GER tasks. We have open-sourced the code and encourage subsequent researchers to broaden the evaluation scope by including more tasks and datasets. Our code and evaluation results are available at: https://github.com/zeroQiaoba/gpt4v-emotion.
keywords:
Generalized Emotion Recognition (GER); GPT-4 with Vision (GPT-4V); Zero-shot Benchmark; Multimodal Fusion; Temporal Modeling.
[…]
[label1] organization=Institute of Automation, Chinese Academy of Sciences, city=Beijing, country=China
[…]
[label2] organization=School of Artificial Intelligence, University of Chinese Academy of Sciences, city=Beijing, country=China
[…]
[label3] organization=Software Engineering Institute, Peking University, city=Beijing, country=China
[…]
[label4] organization=Department of Automation, Tsinghua University, city=Beijing, country=China
[…]
[label5] organization=Beijing National Research Center for Information Science and Technology, Tsinghua University, city=Beijing, country=China
1 Introduction
Emotion has attracted increasing attention from researchers due to its importance in human-computer interaction. Current research on emotion recognition primarily centers on two aspects: one aims to identify the emotions evoked by stimuli, essentially predicting how viewers might feel after seeing these stimuli you2017visual ; the other aims to analyze the emotions conveyed by humans in various ways lian2022smin . We collectively refer to these tasks as “Generalized Emotion Recognition (GER)”.
Emotions are associated with lexical, visual, and acoustic information. Among them, visual information (such as colorfulness, brightness, facial expression, human action, etc) contains rich emotion-related content yang2023emoset . To enhance visual understanding capabilities, researchers have proposed various algorithms and achieved noteworthy progress. With the development of deep learning, current research has shifted from handcrafted features to deep neural networks. Recently, GPT-4V has showcased impressive visual understanding capabilities across various tasks and domains. This leads to a question: can GPT-4V solve the GER problem to some extent? If so, what is the future direction following the emergence of GPT-4V?
In September 2023, GPT-4V was integrated into ChatGPT, triggering a series of user reports aimed at investigating its visual capabilities yang2023dawn . At that time, OpenAI had not yet released the API service, and users could only manually upload test samples to the web service. Due to the high manual effort required, these reports generally involved a limited number of samples per task and merely provided qualitative insights into GPT-4V. In November 2023, OpenAI released the API, but it was initially limited to 100 requests per day. It remained difficult to evaluate GPT-4V against state-of-the-art systems on benchmark datasets. Recently, OpenAI has increased the daily limit, allowing us to conduct more comprehensive evaluations.
In this paper, we provide quantitative evaluation results of GPT-4V on GER tasks, covering visual sentiment analysis ortis2020survey , tweet sentiment analysis niu2016sentiment , micro-expression recognition li2022deep , facial emotion recognition li2020deep , dynamic facial emotion recognition wang2022ferv39k , and multimodal emotion recognition lian2024merbench . Figure 1 shows the overall results of GPT-4V. We also report results of random guessing (a baseline that randomly selects labels from candidate classes) and supervised systems. To fairly compare different methods, we conduct experiments on benchmark datasets and use the same evaluation metric. Overall, GPT-4V outperforms random guessing but still lags behind supervised systems. To figure out the reason, we further conduct a comprehensive analysis of GPT-4V’s multi-faceted capabilities, including multimodal fusion, temporal modeling, robustness, stability, etc. We hope this paper can inspire subsequent researchers and shed light on tasks that GPT-4V can effectively address and those that require further exploration. The main contribution of this paper can be summarized as follows:
  • 1.
    To the best of our knowledge, this is the first work to quantitatively assess GPT-4V’s performance in GER tasks.
  • 2.
    Our evaluation results indicate that GPT-4V is primarily designed for general-purpose domains. It performs poorly in micro-expression recognition that requires specialized knowledge. Meanwhile, it can integrate multimodal clues and capture temporal information during inference.
  • 3.
    This paper can provide guidance on potential future directions following the emergence of GPT-4V. It can also serve as a zero-shot benchmark for subsequent research.
The remainder of this paper is organized as follows: In Section 2, we briefly review recent works. In Sections 3 and 4, we provide an overview of GER tasks and datasets, along with a detailed description of our GPT-4V calling strategy. In Section 5, we report results and conduct an in-depth analysis of GPT-4V’s multi-faceted capabilities. In Sections 6 and 7, we summarize the main challenges and limitations of GPT-4, discuss potential directions for future research, and conclude the entire paper.
Figure 1: Performance of different methods on GER tasks. Here, “random” is a heuristic baseline that randomly selects labels from candidate categories.
2 Related Works
2.1 Generalized Emotion Recognition
GER covers a variety of tasks and this paper focuses on 6 representative tasks. In this section, we detail the similarities and distinctions among them. Visual sentiment analysis ortis2020survey explores the impact of image stimuli on people’s emotions, while tweet sentiment analysis niu2016sentiment aims to determine how individuals feel when posting visual and textual content on social networks. In these tasks, the image is not necessarily human-centric. In contrast, facial emotion recognition li2020deep aims to identify emotions in human-centric images, particularly focusing on observable macro-expressions. Besides macro-expressions, there is another type of expression known as micro-expressions, which are characterized by short duration, low intensity, and sparse facial action units. The task of identifying such expressions is termed micro-expression recognition li2022deep . All of the aforementioned tasks identify emotions from static images. Dynamic facial emotion recognition wang2022ferv39k extends the analysis from static images to image sequences, requiring the use of temporal information. Furthermore, emotions can be conveyed through multiple modalities. Therefore, multimodal emotion recognition lian2024merbench is introduced to integrate multimodal clues for a more comprehensive understanding of emotions.
2.2 Unimodal Emotion Recognition
Emotions can be conveyed through multiple modalities, such as images, text, and audio. Since the early 1970s, Ekman et al. ekman1978facial ; ekman1994strong have conducted prior studies on facial expressions and developed the Facial Action Coding System (FACS) to code facial expressions. In this system, the face can be described as a set of action units, each associated with a muscular basis. Besides facial expressions, emotions can also be conveyed through text and audio, which carry important communicative messages. Text and audio are highly correlated. On the one hand, we can recognize the text from audio. On the other hand, if we ignore the manner in which the text is spoken, we might misunderstand the meaning of the text sebe2005multimodal . With the development of deep learning, researchers have proposed various effective frameworks for emotion recognition lian2021ctnet . These models are usually task-specific, i.e., we need to train on emotion corpora. In contrast, this paper uses GPT-4V, which can recognize emotions in a zero-shot manner without further training.
2.3 Multimodal Emotion Recognition
Humans express emotions in a complementary manner across various modalities. To better understand emotions, we need to integrate multimodal clues. According to the fusion location, current methods can be roughly divided into three categories: feature-level fusion, model-level fusion, and decision-level fusion wu2014survey . Specifically, feature-level fusion concatenates unimodal features at the input level, while decision-level fusion combines unimodal decision results. Despite their promising performance, these methods ignore interactions and correlations between different modalities. Therefore, researchers propose model-level fusion to capture multimodal dependencies. Recently, Lian et al. lian2024merbench built MERBench and conducted a systematic analysis of different model-level fusion strategies. Unlike the above methods, GPT-4V achieves multimodal fusion through prompts. More details can be found in Table 2.
Table 1: Statistics for different datasets. For each dataset, we evaluate performance on the official test set and select the most common evaluation metric. In this table, ACC, WAR, and WAF represent accuracy, weighted average recall, and weighted average F-score, respectively.
Task (Modality) Dataset # Test samples Metric Labels
Visual Sentiment Analysis (Image) Twitter I you2015robust 1,269 ACC positive, negative
Twitter II borth2013large 603 ACC positive, negative
Abstract you2016building 228 ACC amusement, sadness, anger, fear, disgust, awe, content, excitement
ArtPhoto you2016building 806 ACC amusement, sadness, anger, fear, disgust, awe, content, excitement
Tweet Sentiment Analysis (Image, Text) MVSA-Single niu2016sentiment 451 ACC positive, neutral, negative
MVSA-Multiple niu2016sentiment 1,702 ACC positive, neutral, negative
Micro-expression Recognition (Image) CASME yan2013casme 195 ACC happiness, sadness, fear, surprise, disgust, repression, contempt, tense
CASME II yan2014casme 247 ACC happiness, sadness, fear, surprise, disgust, repression, others
SAMM davison2016samm 159 ACC happiness, sadness, fear, surprise, disgust, contempt, anger, others
Facial Emotion Recognition (Image) CK+ lucey2010extended 981 ACC happiness, sadness, anger, fear, disgust, surprise, contempt
SFEW 2.0 dhall2015video 436 ACC happiness, sadness, anger, fear, disgust, surprise, neutral
RAF-DB li2017reliable 3,068 ACC happiness, sadness, anger, fear, disgust, surprise, neutral
FERPlus barsoum2016training 3,589 ACC happiness, sadness, anger, fear, disgust, surprise, neutral, contempt
AffectNet mollahosseini2017affectnet 4,000 ACC happiness, sadness, anger, fear, disgust, surprise, neutral, contempt
Dynamic Facial Emotion Recognition (Video) DFEW (fd1) jiang2020dfew 2,341 WAR happiness, sadness, anger, fear, disgust, surprise, neutral
FERV39k wang2022ferv39k 7,847 WAR happiness, sadness, anger, fear, disgust, surprise, neutral
RAVDESS livingstone2018ryerson 1,440 WAR happiness, sadness, anger, fear, disgust, surprise, neutral, calm
eNTERFACE05 martin2006enterface 1,287 WAR happiness, sadness, anger, fear, disgust, surprise
Multimodal Emotion Recognition (Video, Audio, Text) CMU-MOSI zadeh2017tensor 686 WAF sentiment intensity
CH-SIMS yu2020ch 457 WAF sentiment intensity
MER-MULTI lian2023mer 411 WAF happiness, sadness, anger, surprise, neutral, worry
Figure 2: Examples of different datasets. We pixelate faces due to the sensitive nature of human identity.
2.4 Multimodal Large Language Model
Large language models (LLMs) with massive parameters have attracted increasing attention due to their unprecedented performance. In GER tasks, visual information plays an important role in emotion recognition, but LLMs only support lexical inputs. Therefore, we focus on multimodal LLMs (MLLMs) that support visual inputs. Among all MLLMs, MiniGPT-4 zhu2023minigpt and LLaVA liu2023visual are two representative models. They enhance the visual capability by aligning the frozen visual encoder with LLMs using a simple projection layer, which preserves most of the information encoded in pre-trained models and achieves good performance in visual understanding. This approach is further extended by other works and applies to support more modalities, such as video li2023videochat , audio zhang2023speechgpt , and depth images su2023pandagpt . Currently, GPT-4V is the state-of-the-art MLLM for various tasks alpaca_eval ; lu2024mathvista . However, its performance in emotion recognition has not been fully studied. Hence, this paper evaluates the performance of GPT-4V on GER tasks, providing a valuable zero-shot benchmark for other MLLMs.
2.5 Evaluation on GPT-4V
With the emergence of GPT-4V, a large number of user reports have surfaced to evaluate its visual ability lin2023mm ; wu2023can . Among them, Yang et al. yang2023dawn conducted the most comprehensive assessment, covering coding, captioning, and reasoning abilities. However, they primarily performed qualitative analysis on limited data and did not disclose the performance gap between GPT-4V and supervised systems on benchmark datasets. To bridge this gap, there have been efforts to quantitatively analyze GPT-4V’s mathematical reasoning lu2024mathvista and visual abilities wu2023gpt4vis , but none have specifically explored its emotional understanding capabilities. In this paper, we provide the first quantitative evaluation of GPT-4V on GER tasks. It is worth noting that the emotion is related to humans. However, human identity is treated as sensitive information in GPT-4V, which raises new requirements for the calling strategy. To this end, we also design a calling strategy specifically for GER tasks.
3 Task Description
In this section, we describe each task and dataset in detail. Table 1 summarizes statistics of different datasets. To ensure a fair comparison with supervised systems, we evaluate performance on the official test set and select the most common evaluation metric. Figure 2 shows examples of different datasets, demonstrating the diversity of dataset types. Some datasets are in-the-wild (such as AffectNet), while others are laboratory-controlled (such as CASME and CK+). Meanwhile, there is a disparity in color spaces across distinct datasets. Some datasets use grayscale images (such as CK+), while others use RGB images (such as CASME and AffectNet).
Visual Sentiment Analysis aims to identify the emotions that individuals might evoke from images. We focus on four datasets: Twitter I you2015robust , Twitter II borth2013large , ArtPhoto you2016building , and Abstract you2016building . Specifically, Twitter I you2015robust and Twitter II borth2013large are collected from social websites. Of the two, Twitter I provides raw annotations from 5 Amazon Mechanical Turk workers. Consistent with previous works, we employ three label processing methods, resulting in Twitter I(3), Twitter I(4), and Twitter I(5). Taking Twitter I(4) as an example, it contains all samples where at least 4 workers assign the same label. Different from Twitter I and Twitter II, ArtPhoto you2016building contains artistic photographs from a photo-sharing site, while Abstract you2016building consists of peer-rated abstract paintings. Both datasets are annotated into eight categories. In line with previous works, we reassign these labels into positive and negative classes and report results for the negative/positive classification task.
Tweet Sentiment Analysis aims to determine how people feel when posting visual and textual content on social networks. We conduct experiments on two benchmark datasets: MVSA-Single niu2016sentiment and MVSA-Multiple niu2016sentiment . These datasets contain image-text pairs collected from Twitter, and annotators assign one of three sentiments (positive, negative, and neutral) to the image and text respectively. MVSA-Single contains 5,129 image-text pairs labeled by a single annotator. MVSA-Multiple is an extended version of MVSA-Single with more samples and more annotators. For a fair comparison, we adopt the same dataset processing strategy in previous works, and then randomly select 10% of the samples as the test set xu2017multisentinet ; zhu2022multimodal .
Facial Emotion Recognition focuses on five benchmark datasets: CK+ lucey2010extended , FERPlus barsoum2016training , SFEW 2.0 dhall2015video , RAF-DB li2017reliable , and AffectNet mollahosseini2017affectnet . The first two datasets contain grayscale images, while the last three datasets contain RGB images. Specifically, CK+ lucey2010extended includes 593 video sequences from 123 subjects. Following previous work jiang2022disentangling , we extract the last three frames of each sequence to construct the dataset. RAF-DB li2017reliable contains thousands of samples with basic and compound expressions, and this paper focuses on basic emotions. AffectNet mollahosseini2017affectnet has 8 labels and each label contains 500 samples for evaluation. For FERPlus barsoum2016training and SFEW 2.0 dhall2015video , we use the official test set for performance evaluation.
Table 2: Prompt template for different tasks. Please replace “Candidate labels” in the following prompts into candidate labels of each dataset. This paper uses batch-wise calling strategy to handle request limits. By default, we set the batch size to 20 for image-level inputs and 6 for video-level inputs.
Task
Prompt
Visual Sentiment Analysis
Please play the role of an emotion recognition expert. We provide 20 images. Please recognize sentiments evoked by these images (i.e., guess how viewer might emotionally feel after seeing these images.) If there is a person in the image, ignore that person’s identity. For each image, please sort the provided categories from high to low according to the similarity with the image. Here are the optional categories: {Candidate labels}. The output format should be {’name’:, ’result’:} for each image.
Tweet Sentiment Analysis
Please play the role of an emotion recognition expert. We provide 20 image-text pairs. Please analyze how he will feel if he post this image-text pair on social media. If there is a person in the image, ignore that person’s identity. For each image-text pair, please sort the provided categories from high to low according to the similarity with the input image-text pair. Here are the optional categories: {Candidate labels}. The output format should be {’name’:, ’result’:} for each image-text pair.
Micro-expression Rec
Please play the role of a micro-expression recognition expert. We provide 20 images. For each image, please sort the provided categories from high to low according to the similarity with the input image. The expression may not be obvious, please pay attention to the details of the face. Here are the optional categories: {Candidate labels}. Please ignore the speaker’s identity and focus on the facial expression. The output format should be {’name’:, ’result’:} for each image.
Facial Emotion Rec.
Please play the role of a facial expression classification expert. We provide 20 images. For each image, please sort the provided categories from high to low according to the top 5 similarity with the input image. Here are the optional categories: {Candidate labels}. Please ignore the speaker’s identity and focus on the facial expression. The output format should be {’name’:, ’result’:} for each image.
Dynamic Facial Emotion Rec.
Please play the role of a video expression classification expert. We provide 6 videos, each with 3 temporally uniformly sampled frames. For each video, please sort the provided categories from high to low according to the top 5 similarity with the input video. Here are the optional categories: {Candidate labels}. Please ignore the speaker’s identity and focus on the facial expression. The output format should be {’name’:, ’result’:} for each video.
Multimodal Emotion Rec.
Please play the role of a video expression classification expert. We provide 6 videos, each with the speaker’s content and 3 temporally uniformly sampled frames. Please ignore the speaker’s identity and focus on their emotions. For each video, please sort the provided categories from high to low according to the top 5 similarity with the input video. Here are the optional categories: {Candidate labels}. Please ignore the speaker’s identity and focus on their emotions. The output format should be {’name’:, ’result’:} for each video.
Micro-expression Recognition aims to identify subtle changes in human faces. Early works have proved the feasibility of recognizing micro-expressions from an apex frame li2020joint . Therefore, we use the apex frame in our evaluation. Following previous works, we adopt the same label processing method and concentrate on major emotions. Specifically, CASME yan2013casme comprises 195 samples across 8 categories, and we focus on the performance of four main labels: tense, disgust, repression, and surprise. CASME II yan2014casme includes 247 samples collected from 26 subjects, and we focus on five major labels: happiness, surprise, disgust, repression, and others. SAMM davison2016samm consists of 159 samples, and our evaluation is limited to labels with more than 10 samples: anger, contempt, happiness, surprise, others.
Dynamic Facial Emotion Recognition focuses on more challenging image sequences. In this task, we conduct experiments on four benchmark datasets: FERV39k wang2022ferv39k , RAVDESS livingstone2018ryerson , eNTERFACE05 martin2006enterface , and DFEW jiang2020dfew . For the first three datasets, we evaluate the performance on the official test set. DFEW jiang2020dfew contains 5 folds with a total of 11,697 samples. To reduce the evaluation cost, we only report the results of fold 1 (fd1), consistent with previous works wang2022dpcnet ; chen2023static .
Multimodal Emotion Recognition focuses on three benchmark datasets: CH-SIMS zadeh2017tensor , CMU-MOSI yu2020ch , and MER-MULTI lian2023mer . The first two datasets offer sentiment intensity scores. In this paper, we concentrate on the negative/positive classification task, where positive and negative classes are assigned for $<0$ and $>0$ scores, respectively. MER-MULTI is a subset of the MER2023 dataset lian2023mer , providing both discrete and dimensional labels for each sample. In this paper, we focus on the recognition performance of discrete emotions.
Input: Dataset $\mathcal{D}=\{x_{i}\}_{i=1}^{N}$ with ${N}$ samples, the maximum number of calls $T_{\max}$.
Output: Prediction results for all samples.
1 Def Main($\mathcal{D}$):
2       Split $\mathcal{D}$ into batches $\mathcal{D}=\{\mathcal{B}_{i}\}_{i=1}^{N_{b}}$; for $i=1,\cdots,N_{b}$ do
3             # batch-wise calling Func($\mathcal{B}_{i}$);
4       end for
5      return
6
7
8 Def Func($\mathcal{B}_{i}$):
9       # repeated calling for $t=1,\cdots,T_{\max}$ do
10             Call GPT-4V for $\mathcal{B}_{i}$ and get response $\mathcal{O}_{i}$; if meet requirements then
11                   Store the response $\mathcal{O}_{i}$; return
12             end if
13
14       end for
15      # recursive calling if $|\mathcal{B}_{i}|\geq 2$ then
16             Divide $\mathcal{B}_{i}$ into two equal parts: $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$; Func($\mathcal{C}_{1}$); Func($\mathcal{C}_{2}$);
17       end if
18      return
19
Algorithm 1 GPT-4V Calling Strategy
4 GPT-4V Calling Strategy
This paper evaluates the performance of the GPT-4V API, “gpt-4-vision-preview”. GER tasks involve diverse modalities, such as images, text, video, and audio. However, GPT-4V only supports image and text inputs. To process video data, we sample the video and convert it into multiple images. To process audio data, we try converting the audio to a mel-spectrogram, which is a typical visual representation of the frequency content of an audio signal over time, with the frequency axis scaled according to the mel scale hermansky1990perceptual . However, GPT-4V fails to generate proper responses for the mel-spectrogram. Therefore, the primary focus of this paper revolves around images, text, and video. In this section, we design a calling strategy specifically for GER tasks, containing batch-wise, repeated, and recursive calling. Pseudo-code is summarized in Algorithm 1.
Batch-wise Calling. GPT-4V API has three limitations on requests: tokens per minute (TPM), requests per minute (RPM), and requests per day (RPD), which introduces additional requirements to our prompt design. To meet the constraints of RPM and RPD, we follow previous work wu2023gpt4vis and adopt batch-wise inputs. That is, we feed multiple samples into GPT-4V and generate all results in one request. However, a large batch size may result in the total number of tokens exceeding the TPM limitation. Additionally, it increases task difficulty and may lead to incorrect responses. For example, if we input 100 samples in a request, GPT-4V might only generate predictions for 96 samples. Therefore, we set the batch size to 20 for image-level inputs and 6 for video-level inputs, aiming to satisfy TPM, RPM, and RPD limitations simultaneously. Prompts for different tasks can be found in Table 2.
Repeated Calling. GER tasks often trigger security checks, causing GPT-4V to refuse to provide a response. This is attributed to the inclusion of visual sentiment analysis and human emotion recognition. The former task contains violent and bloody images. In the latter task, human identities are also considered as sensitive information. To reduce rejected cases, we require GPT-4V to ignore speaker identity (see Table 2), but it still triggers security errors. Interestingly, these errors sometimes occur randomly. For example, although all images are human-centric, some pass security checks while others fail. Alternatively, a sample may initially fail the check but pass upon retry. Therefore, we make multiple calls to the rejected batches.
Recursive Calling. During the evaluation, we notice that a batch-wise input may fail security checks, but splitting it into smaller parts can sometimes pass checks. Therefore, for the consistently failing batch, we split it into two smaller mini-batches and feed them into GPT-4V separately. This operation is repeated until further splitting is no longer feasible.
Combination. Our strategy combines batch-wise, repeated, and recursive calling. More details can be found in Algorithm 1. A qualified response $\mathcal{O}$ should satisfy two conditions. First, it should not trigger the security check. Second, it should contain the correct number of prediction results. Figures 8$\sim$13 provide examples of qualified responses for different GER tasks.
5 Results and Discussion
This section presents and discusses results from three levels: dataset-wise, class-wise, and sample-wise. Specifically, we first present the results of heuristics baselines, supervised systems, and GPT-4V. Then, we reveal the temporal modeling and multimodal fusion capabilities of GPT-4V. Following that, we analyze the system’s stability and conduct an in-depth class-wise performance analysis. Next, we assess the system’s robustness against changes in prompt templates and color spaces. Finally, we visualize the rejected and erroneously predicted samples.
Table 3: Performance comparison on visual sentiment analysis.
Model Twitter I (5/4/3) Twitter II Abstract ArtPhoto
SentiBank borth2013large 71.32/68.28/66.63 65.93 64.95 67.74
PAEF zhao2014exploring 72.90/69.61/67.92 77.51 70.05 67.85
DeepSentiBank chen2014deepsentibank 76.35/70.15/71.25 70.23 71.19 68.73
PCNN you2015robust 82.54/76.52/76.36 77.68 70.84 70.96
VGGNet simonyan2015very 83.44/78.67/75.49 71.79 68.86 67.61
AR (Concat) yang2018visual 88.65/85.10/81.06 80.48 76.03 74.80
Random 50.74/49.75/50.50 50.37 50.88 50.41
Majority 66.82/62.51/61.33 78.61 61.23 53.06
GPT-4V 97.81/94.63/90.71 87.95 71.81 80.40
Table 4: Performance comparison on micro-expression recognition.
Model CASME CASME II SAMM
LBP-SIP wang2015lbp 36.84 46.56 36.76
STRCN-A xia2019spatiotemporal 40.93 45.26 32.85
STRCN-G xia2019spatiotemporal 59.65 63.37 53.48
VGGMag li2018can 60.23 63.21 36.00
LGCcon li2020joint 60.82 65.02 40.90
TSCNN song2019recognizing 73.88 80.97 71.76
Random 24.09 18.33 20.37
Majority 40.34 40.17 42.22
GPT-4V 36.93 14.64 17.04
Table 5: Performance comparison on tweet sentiment analysis.
Model MVSA-Single MVSA-Multiple
CNN-Multi cai2015convolutional 61.20 66.39
DNN-LR yu2016visual 61.42 67.86
MultiSentiNet xu2017multisentinet 69.84 68.86
CoMN xu2018co 70.51 69.92
MVAN yang2020image 72.98 72.36
ITIN zhu2022multimodal 75.19 73.52
Random 33.96 33.46
Majority 55.68 66.65
GPT-4V 61.25 66.82
Table 6: Performance comparison on multimodal emotion recognition. These results come from MERBench lian2024merbench , which reproduces different algorithms using the same feature set to ensure a fair comparison.
Model CMU-MOSI CH-SIMS MER-MULTI
MFM tsai2018learning 78.34 87.14 70.48
MISA hazarika2020misa 79.08 89.82 82.42
MFN zadeh2018memory 77.78 87.00 77.40
MMIM han2021improving 79.38 88.68 81.01
TFN zadeh2017tensor 79.63 90.56 82.52
MulT tsai2019multimodal 81.41 91.07 82.94
Random 51.33 51.23 17.87
Majority 42.37 50.47 10.40
GPT-4V 80.43 81.24 65.39
Table 7: Performance comparison on facial emotion recognition.
Model CK+ Model SFEW 2.0 Model RAF-DB Model FERPlus Model AffectNet
Cross-VAE wu2020cross 94.96 IACNN meng2017identity 50.98 SCN wang2020suppressing 87.03 RANwang2020region 89.16 SCN wang2020suppressing 60.23
IA-gen yang2018identity 96.57 DLP-CNN li2017reliable 51.05 EfficientFace zhao2021robust 88.36 SCN wang2020suppressing 89.39 MA-Net zhao2021learning 60.29
ADFL bai2019disentangled 98.17 TDTLN yan2019cross 53.10 MA-Net zhao2021learning 88.42 EAC zhang2022learn 89.64 ARM shi2021learning 61.33
IDFERM liu2019hard 98.35 DAN wen2023distract 53.18 TransFER xue2021transfer 90.91 KTNli2021adaptively 90.49 DAN wen2023distract 62.09
TER-GAN ali2021facial 98.47 RAN wang2020region 54.19 POSTER zheng2023poster 92.05 TransFERxue2021transfer 90.83 POSTER zheng2023poster 63.34
IPD-FER jiang2022disentangling 98.65 FaceCaps wu2021facecaps 58.50 POSTER++ mao2023poster 92.21 POSTER zheng2023poster 91.62 POSTER++ mao2023poster 63.77
Random 14.61 Random 13.98 Random 14.57 Random 12.61 Random 12.73
Majority 25.38 Majority 19.77 Majority 38.64 Majority 35.75 Majority 12.50
GPT-4V 69.72 GPT-4V 57.24 GPT-4V 75.81 GPT-4V 64.25 GPT-4V 42.77
Table 8: Performance comparison on dynamic facial emotion recognition. Since GPT-4V does not support video input directly, we uniformly sample frames from the videos and sequentially feed these frames into GPT-4V. In this table, we test 2-frame (2-frm) and 3-frame (3-frm) sampling manners.
Model DFEW Model FERV39k Model RAVDESS Model eNTERFACE05
C3D tran2015learning 56.37 M3DFEL wang2023rethinking 47.67 AV-LSTM ghaleb2019multimodal 65.80 STA-FER pan2019deep 42.98
ResNet50-LSTM wang2022dpcnet 63.32 STT ma2022spatio 48.11 MCBP su2020msaf 71.32 EC-LSTM miyoshi2021enhanced 49.26
DPCNet wang2022dpcnet 65.78 IAL li2023intensity 48.54 MSAF su2020msaf 74.86 FAN meng2019frame 51.44
SVFAP sun2023svfap 74.81 MAE-DFER sun2023mae 52.07 SVFAP sun2023svfap 75.01 Graph-Tran zhao2022spatial 54.62
MAE-DFER sun2023mae 74.88 SVFAP sun2023svfap 52.29 MAE-DFER sun2023mae 75.56 SVFAP sun2023svfap 60.54
S2D chen2023static 76.16 S2D chen2023static 52.56 CFN-SR fu2021cross 75.76 MAE-DFER sun2023mae 61.64
Random 14.43 Random 14.41 Random 11.86 Random 16.87
Majority 22.81 Majority 24.79 Majority 13.33 Majority 16.71
GPT-4V (2-frm) 43.80 GPT-4V (2-frm) 34.29 GPT-4V (2-frm) 22.29 GPT-4V (2-frm) 23.78
GPT-4V (3-frm) 54.80 GPT-4V (3-frm) 38.72 GPT-4V (3-frm) 34.31 GPT-4V (3-frm) 33.28
5.1 Main Results
Besides GPT-4V, we introduce two heuristic baselines: random guessing and majority guessing. For the former, we randomly select a label from candidate labels. For the latter, we choose the majority label as the prediction. For both baselines, we run experiments 10 times and report average results.
Table 3 shows the results of visual sentiment analysis. We observe that GPT-4V outperforms supervised systems on most datasets. This superior performance can be attributed to its strong visual understanding ability, coupled with its reasoning capability, allowing GPT-4V to accurately infer the emotional states evoked by images. But for micro-expression recognition (see Table 4), GPT-4V exhibits poor performance, even worse than heuristic baselines. These results suggest that GPT-4V is designed for general-purpose domains (i.e., emotions that can be perceived by normal people). It is not well-suited for tasks that require professional knowledge like micro-expressions.
Tables 5$\sim$8 present the results of tweet sentiment analysis, multimodal emotion recognition, facial emotion recognition, and dynamic facial emotion recognition, respectively. To process video data, we uniformly sample frames and feed these frames into GPT-4V sequentially. To reduce call costs, we sample up to three frames per video. Despite a performance gap still existing between GPT-4V and supervised systems, it is worth noting that GPT-4V can significantly outperform heuristic baselines, demonstrating its potential in emotion recognition.
In Table 6, we observe a larger performance gap between supervised systems and GPT-4V on MER-MULTI compared to CMU-MOSI and CH-SIMS. This can be attributed to the fact that supervised systems incorporate acoustic information, while GPT-4V does not support audio input. Given that audio is more crucial in MER-MULTI lian2024merbench , GPT-4V loses more information in this dataset, leading to limited performance.
5.2 Temporal Modeling Ability
To reduce GPT-4V call costs, this paper samples up to three frames per video. In this section, we further reveal the impact of sampling numbers. In Table 8, the performance improves when we increase the sampling number from two to three, indicating GPT-4V’s temporal understanding ability. Additionally, it is worth noting that despite setting the sampling number to three, some key frames may still be ignored. Therefore, there is a possibility for further improvement by sampling more frames, and we leave this discussion for subsequent researchers.
5.3 Multimodal Fusion Ability
This section evaluates GPT-4V’s multimodal fusion ability. Among all tasks, tweet emotion recognition and multimodal emotion recognition provide multimodal information. Hence, we conduct experiments on these tasks. Table 9 shows unimodal and multimodal results. In general, multimodal results outperform unimodal results, showcasing GPT-4V’s ability to integrate and exploit multi-modalities. But for CMU-MOSI, we notice a slight decrease in multimodal results compared to unimodal results. This can be attributed to CMU-MOSI mainly relying on lexical content to convey emotions lian2024merbench , and the incorporation of visual clues may introduce interference.
Table 9: Multimodal fusion ability. We report unimodal and multimodal results for tweet sentiment analysis (TSA) and multimodal emotion recognition (MER). Here, TSA takes text-image inputs and MER takes text-video inputs.
Task Dataset Text Image/Video Fusion
TSA MVSA-Single 57.65 58.68 61.25
MVSA-Multiple 62.71 63.35 66.82
MER CH-SIMS 70.07 76.13 81.24
CMU-MOSI 82.32 51.17 80.43
MER-MULTI 34.57 46.23 65.39
5.4 System Stability
This section evaluates GPT-4V’s prediction stability. During the experiments, we run GPT-4V 10 times for each sample in SFEW 2.0. Figure 3a shows the frequency of identical predictions. Specifically, we assume that GPT-4V predicts negative 8 times and negative 2 times. Therefore, it predicts the same label at most $c=8$ times. Then, we compute $c$ for all samples and calculate its frequency. In Figure 3b, we report the test accuracy for each run. We observe that although more than 50% of the samples exhibit the same result across 10 runs, there are also some samples with different prediction results across distinct runs, leading to fluctuations in test accuracy. Meanwhile, we observe a 4.60% performance gap between the best and worst runs. Therefore, GPT-4V exhibits certain instability. We recommend that subsequent researchers evaluate GPT-4V multiple times and use majority voting to get final results.
5.5 Class-wise Performance Analysis
In Figure 4, we visualize confusion matrices and perform a class-wise performance analysis. For visual sentiment analysis, GPT-4V generally performs well in positive and negative classes, showcasing its visual understanding ability. However, it exhibits lower results on Abstract compared to other datasets. This may be attributed to the fact that GPT-4V is mainly trained on natural images. The domain gap between abstract and natural images leads to increased difficulty and limited performance. For tweet sentiment analysis, GPT-4V performs relatively poorly in identifying neutral sentiment and often misclassifies it as positive or negative. This may be due to GPT-4V’s ability to detect subtle emotional cues in image-text pairs.
(a) identical predictions
(b) test accuracy
Figure 3: System stability test. We run GPT-4V 10 times, calculate the frequency with identical predictions, and report the test accuracy for different runs.
(a) Twitter I(5)
(b) Twitter II
(c) Abstract
(d) ArtPhoto
(e) MVSA-Single
(f) MVSA-Multiple
(g) CASME
(h) CASME II
(i) SAMM
(j) CK+
(k) SFEW 2.0
(l) FERPlus
(m) RAF-DB
(n) AffectNet
(o) DFEW
(p) FERV39k
(q) RAVDESS
(r) eNTERFACE05
(s) CMU-MOSI
(t) CH-SIMS
(u) MER-MULTI
Figure 4: Confusion matrices for all evaluation datasets. Different colors represent distinct tasks.
Table 10: Robustness to prompt template changes. We conduct experiments on the facial emotion recognition task.
ID
Prompt Template
#1
Please play the role of a facial expression classification expert. We provide 20 images. For each image, please sort the provided categories from high to low according to the top 5 similarity with the input image. Here are the optional categories: {Candidate labels}. Please ignore the speaker’s identity and focus on the facial expression. The output format should be {’name’:, ’result’:} for each image.
#2
We provide 20 images. For each image, please sort the provided categories from high to low according to the top 5 similarity with the input image. Here are the optional categories: {Candidate labels}. Please ignore the speaker’s identity and focus on the facial expression. The output format should be {’name’:, ’result’:} for each image.
#3
Please play the role of a facial expression classification expert. We provide 20 images. For each image, please select the most likely category according to the correlation with the input image. Here are the optional categories: {Candidate labels}. Please ignore the speaker’s identity and focus on the facial expression. The output format should be {’name’:, ’result’:} for each image.
For micro-expression recognition, GPT-4V prefers tense and repression. Compared to other labels (e.g., surprise and disgust), tense and repression are less noticeable and closer to neutral. These results indicate that GPT-4V tends to view micro-expressions as neutral, causing it to perform poorly on this task. For facial expression recognition, GPT-4V tends to exhibit lower accuracy on contempt, disgust, and fear, while achieving higher accuracy on happiness. This phenomenon is related to the clarity of emotion definition. Since contempt, disgust, and fear have less consensus between different annotations, they are inherently more difficult to recognize mollahosseini2017affectnet ; picard2001toward .
For dynamic facial emotion recognition, most samples tend to be predicted as neutral, which is more obvious in FERV39. We attribute this phenomenon to our sampling operation. For each video, we sample up to three frames. GPT-4V tends to predict the emotion as neutral if there is no emotion-related information within these frames. For multimodal emotion recognition, GPT-4V performs well on worry and happiness. For worry, it has a strong reliance on lexical information. Differently, happiness is mainly expressed through facial expressions lian2023mer . To achieve good performance, GPT-4V should use different predominant modalities for distinct emotions, showcasing its multimodal fusion capabilities.
5.6 Robustness to Template Change
Previous works have pointed out that pre-trained language models (such as BERT and RoBERTa) are sensitive to prompt templates, with different templates yield varying results mao2022biases . This raises the question: is GPT-4V also sensitive to template changes? In this section, we conduct experiments on SFEW 2.0 and compare the performance of different templates. In Table 10, Template#1 is our default template. Compared with Template#1, Template#2 deletes “Please play the role of a facial expression classification expert” and Template#3 changes the selection of the top 5 classes into the selection of the most likely class. Experimental results demonstrate that Template#2 (57.93%) and Template#3 (54.25%) can achieve similar performance compared to Template#1 (57.24%). Therefore, we can conclude that GPT-4V is robust to template changes.
5.7 Robustness to Color Space
This section evaluates the robustness of GPT-4V to color space changes. We conduct experiments on RAF-DB and convert RGB images into grayscale images. We observe that GPT-4V achieves nearly identical results on grayscale (74.28%) and RGB images (75.81%). These results demonstrate its robustness to color space changes. In Figure 5, we further analyze the class-wise prediction consistency. We observe that the consistency rate is around 80% for most classes except for fear and disgust. We further compare Figure 5 with the confusion matrix in Figure 4m. We observe similarities between these figures, where fear and disgust also perform poorly in emotion recognition. Therefore, we can conclude that poorly-performing categories in GPT-4V exhibit greater inconsistency in different color spaces. This may be attributed to poorly-performing categories with greater randomness across different runs.
Figure 5: Prediction consistency on RGB and grayscale images.
Figure 6: Passed and rejected examples by GPT-4V.
5.8 Security Check
GER includes visual sentiment analysis (containing images with violent and bloody content) and human emotion recognition (where human identity is considered sensitive information in GPT-4V). Therefore, it often triggers the security check. Figure 6 visualizes passed and rejected examples. In general, rejected examples contain humans or violent scenes. But there are some cases where examples containing such information can still pass checks. This observation suggests a certain degree of instability in the current security check used by GPT-4V.
5.9 Case Study
In Figure 7, we visualize examples where GPT-4V makes incorrect predictions. More examples can be found in Figures 8$\sim$13 in the Appendix. For visual sentiment analysis (see Twitter I), we notice that GPT-4V can generate reasonable predictions. The main reason for these errors is the inherent multiple meanings of images. Take the image on the left as an example. One interpretation could emphasize that the sun rises, breaking through dark clouds and bathing the trees in sunlight, evoking positive emotions. Conversely, another explanation could emphasize that the sun is obstructed by dark clouds, triggering negative emotions. The same phenomenon can also be found in the picture on the right. On the one hand, there is a cute cat, but on the other hand, the cat is locked in a cage.
Figure 7: Error cases predicted by GPT-4V.
For tweet sentiment analysis (see MVSA-Multiple), the image shows a hat on wood, accompanied by descriptive text about the hat. Although the label is neutral, GPT-4V predicts a positive sentiment. We speculate that GPT-4V may interpret this image as depicting a comfortable environment and a nice hat, thereby eliciting a positive response.
For micro-expression recognition (see CASME), we observe that her emotion is not obvious. Without specialized knowledge, it is difficult to identify her emotion as disgust and more likely to be predicted as neutral. For CASME, there are four candidate labels: tense, disgust, repression, and surprise. Among them, tense is closer to neutral. Consequently, GPT-4V incorrectly predicts her emotion as tense. In the future, we recommend that researchers provide clearer definitions of micro-expressions in the prompt. For example, offer typical clues to distinguish these micro-expressions. We believe this strategy can enhance the performance of GPT-4V on this task.
For facial emotion recognition (see RAF-DB), the image depicts a baby with an open mouth, indicating that she is crying. Meanwhile, she is wearing an oxygen mask, indicating that she is sick. Hence, her emotion is likely to be sad. But GPT-4V incorrectly predicts her emotion as neutral. This may be due to its inability to recognize the oxygen mask and infer her health status, indicating its limitations in understanding the details.
For dynamic facial emotion recognition (see DFEW), GPT-4V might predict the emotion as happiness based on the raised corners of the mouth. But when considering the subtitle, “cannot go to the academy, read books, and listen to music”, we can infer that he is more likely to be sad due to these restrictions. Hence, the error made by GPT-4V could be attributed to the limitations of this task, as it relies only on the visual modality.
For multimodal emotion recognition (see CH-SIMS), GPT-4V predicts the emotion as positive based on his facial expression. However, when taking into account the subtitle, “I feel like I’m falling apart”, the content of his complaint shifts his emotion towards negative. However, GPT-4V is unable to capture such emotional transitions, indicating its limitations in multimodal fusion under complex scenarios.
In summary, GPT-4V can provide relatively reasonable responses. These errors are mainly caused by the inherent multiple meanings of images or the limited capability to understand the details. Meanwhile, micro-expressions are different from regular emotions and GPT-4V faces challenges in recognizing them. Furthermore, its multimodal fusion ability may face challenges when different modalities express distinct emotions.
6 Limitations and Future Directions
In this section, we summarize the main challenges and limitations of GPT-4V as revealed by the evaluation results. Additionally, we discuss potential directions for future research.
Performance. In GER tasks, GPT-4V outperforms heuristic baselines but still lags behind supervised systems. This may be attributed to the vague definition of emotion lian2023explainable . In the future, we plan to use few-shot prompts to enhance task clarity and GPT-4V’s emotion understanding ability.
Support Modality. GPT-4V does not support audio. However, audio can convey emotions through prosodic information such as pitch, duration, and intensity. To better recognize emotions, GPT-4V should support audio in the future. Alternatively, we can use outside models to convert audio into lexical descriptions and use them as additional input for GPT-4V.
Domain Gap. GPT-4V is designed for general-purpose domains and performs poorly on micro-expressions that require specialized knowledge. For future work, we recommend that subsequent researchers can attempt to provide clearer definitions of micro-expressions in the prompt design.
System Stability. GPT-4V exhibits certain instability in prediction results, i.e., it may predict different labels for a sample. Therefore, GPT-4V should improve its system stability in the future. Alternatively, subsequent works should evaluate GPT-4V multiple times and use majority voting to get final results.
Security Check Stability. GPT-4V demonstrates certain instability in security checks. Specifically, a sample may initially fail the check but pass after a retry. To solve this problem, we design a more practical calling strategy in Section 4. In the future, GPT-4V should enhance its security check stability.
Despite the above limitations, GPT-4V has demonstrated promising results in emotion recognition and holds application potential in practical scenarios such as social media analysis, online education, and customer service. Different from previous solutions that required task-specific models tailored to each scenario, GPT-4V can serve as a universal model capable of minimizing application costs.
7 Conclusion
This paper conducts a comprehensive evaluation of GPT-4V in GER tasks. Our evaluation results demonstrate that GPT-4V exhibits strong visual understanding capabilities and achieves promising results in general-purpose emotion recognition. However, it performs poorly in micro-expression recognition which requires specialized knowledge. This paper also showcases GPT-4V’s temporal modeling and multimodal fusion capabilities, along with its robustness to color space and prompt template changes. Furthermore, we evaluate its prediction consistency and security check stability and visualize error cases to reveal its limitations in emotion understanding. This paper can serve as a zero-shot benchmark and provide guidance for subsequent research on emotion recognition and MLLMs.
Acknowledgements
This work is supported by the National Natural Science Foundation of China (NSFC) (No.62201572, No.61831022, No.62276259, No.62322120, No.U21B2010) and Beijing Municipal Science & Technology Commission, Administrative Commission of Zhongguancun Science Park (No.Z211100004821013).